OpenStudy (jennychan12):

What is this asking: Graph f(x) = cosx-sinx. Calculate the area in the first quadrant bounded by the x-axis, the y-axis and f?

OpenStudy (jennychan12):

that's the whole question, but i'm not given the limits or anything.

OpenStudy (anonymous):

I think your condition is x>0 and y>0, and the boundaries is x,y and f(x)

OpenStudy (jennychan12):

ok, but what are the limits that i have to integrate?

OpenStudy (campbell_st):

I think there needs to be a domain for x the lower boundary is 0 and whats the upper boundary.. 2pi

OpenStudy (anonymous):

graph it, I think the integration is from x=0 to where f(x) = 0, but as the way I graphed it here http://my.hrw.com/math06_07/nsmedia/tools/Graph_Calculator/graphCalc.html Doesn't look like a finite answer if there is no stop point

OpenStudy (jennychan12):

yeah that's what i'm confused about because it seems to be an indefinte integral...

OpenStudy (campbell_st):

well if its 1st quadrant the domain is (0, pi/4)

OpenStudy (campbell_st):

let y = 0 so solve cos(x) - sin(x) = 0 or cos(x) = sin(x) this only occurs when x = pi/4

OpenStudy (jennychan12):

ok thanks. cuz the answer is less than 1

OpenStudy (anonymous):

seem like it's the first portion of the curve on the first quadrant, how about the rest?

OpenStudy (campbell_st):

but 1st quadrant may require you to integrate between (0 and pi/2)

OpenStudy (jennychan12):

ok thanks :D

OpenStudy (campbell_st):

so you will need to split it into 2 integrals as pi/4 to pi/2 is below the x-axis so could use \[\int\limits_{0}^{\frac{\pi}{4}} \cos(x) - \sin(x) dx + \int\limits_{\frac{\pi}{4}}^{\frac{\pi}{2}} \cos(x) - \sin(x) dx \] or \[A = 2\int\limits_{0}^{\frac{\pi}{4}} \cos(x) - \sin(x) dx\]